The lower base angles of an isosceles trapezoid are labeled a and b. What is the size of the angles if a is equal to 5x + 3 and b is equal to 3x + 17?
Answer by ewatrrr(24785)

 
Hi,
The lower base angles of an isosceles trapezoid are labeled a and b.
What is the size of the angles if a is equal to 5x + 3 and b is equal to 3x + 17
The base angles of an isosceles trapezoid are congruent(same)
5x+3 = 3x+17
2x = 14
x = 7, 5x + 3 = 38° and 3x+17 = 38°